a. what is the probability that a couple (with no children) will have three consecutive female children? b. What is the probability that a couple with 2 female children will have another female child as a result of their next pregnancy?

please type the answer.
Answer #1

a) having a female child is independent of previous pregnancies

female is XX and male XY

XY* XX

X X
X XX ( female) XX ( female)
Y XY ( male) XY ( male)

Probability of having a female child= number of females/ total number

=2/4

=1/2

so the probability of having 3 female children =( Probability of having a female child )^3

= (1/2)^3

= ( 1/8)

b) having a child is an independent event so probability that a couple with 2 female children will have another female child as a result of their next pregnancy= probability of getting a female child= 1/2